OpenStudy (anonymous):

Solve the equation for x: 2x-16y=10 please explain? Thank you. (:

OpenStudy (anonymous):

My guess is y=32 and x=......Nevermind :P

OpenStudy (anonymous):

Is there another equation with it?

OpenStudy (anonymous):

so first of all you can divide straight across by 2 x+8y=5 so add 8y to both sides x-8y+8y=5+8y x=5+8y

OpenStudy (anonymous):

Hi Welcome to Openstudy! @Christina_212 @DanZander @CynfulBaddie @TRavi14 Are you looking for the x-intercept? or as @TRavi14 asked, do you have any other equations?

OpenStudy (anonymous):

@CynfulBaddie I do not think you did it correctly. If it is done like that you must subtract 8y from both sides not add.

OpenStudy (anonymous):

2x - 16y = 10 Add 16y to both sides 2x - 16y + 16y = 16y + 10 2x = 16y + 10 Divide everything by 2 x = 8y+5

OpenStudy (anonymous):

Yes!